Answer the next question on the basis of the following information about a hypothetical economy.

Full-time employed = 80

Part-time employed = 25

Unemployed = 15

Discouraged workers = 5

Members of underground economy = 6

Consumer price index = 110

The unemployment rate is
Answer #1

Labor force = 80 + 25 + 15 = 120

Unemployed = 15

Unemployment rate = unemployed / labor force * 100

= 15 / 120 * 100

= 12.50%